OpenStudy (anonymous):

Question 4 of 12 The formula for the perimeter of a rectangle is P = 2l + 2w. The length of a rectangle is 7 units longer than its width. Which expression represents the perimeter of the rectangle? 2 + w(7w) 2 + w(w + 7) 2(7w) + 2w 2(w + 7) + 2w

OpenStudy (anonymous):

i think its b

OpenStudy (danjs):

welcome to open study flower

OpenStudy (anonymous):

thank you

OpenStudy (danjs):

Lets take the information 1 sentence at a time ok

OpenStudy (danjs):

Perimeter for a rectangle is: P = 2*L + 2*W

OpenStudy (danjs):

"The length is 7 units longer than the width" L = W + 7

OpenStudy (anonymous):

i think this answer is b im kinda not sure though

OpenStudy (danjs):

we will get there... you see how i got L = W + 7 from " the length is 7 units longer than the width?

OpenStudy (anonymous):

yes

OpenStudy (danjs):

Perimeter: P = 2*L + 2*W Now we are told : L = W + 7, put that into the perimeter above for L P = 2*(W+7) +2*W

OpenStudy (danjs):

You see how i just replaced the L in the perimeter , with the info they told you , the length L is 7 larger than the width W. L= W+ 7

OpenStudy (anonymous):

oh wow that made alot of since thank you

OpenStudy (danjs):

no problem, anytime, If you ask a question you can type @DanJS and it will get my attention

OpenStudy (anonymous):

ok great! :)

OpenStudy (anonymous):

do i close this now?

OpenStudy (danjs):

If you would like to add me , you can hover over my name and click the Become a Fan button,

OpenStudy (danjs):

yeah , close the question

OpenStudy (anonymous):

ok i did fan you bye

OpenStudy (danjs):

thanks